OpenStudy (anonymous):

WILL FAN & MEDAL!!!! please help!! For this question, use integer values of x from -3 to 3 to graph the equation. y=-2x^2+3

OpenStudy (anonymous):

a.

OpenStudy (anonymous):

b.

OpenStudy (anonymous):

c.

OpenStudy (anonymous):

d.

OpenStudy (is3535):

Step 1 : Raise x to the 2nd power Exponentiaion : Equation at the end of step 1 : y - ((2 • x2) + 3) = 0 Step 2 : Simplify y - 2x2+3 Trying to factor a multi variable polynomial : 2.1 Factoring y - 2x2 - 3 Try to factor this multi-variable trinomial using trial and error Factorization fails Equation at the end of step 2 : y - 2x2 - 3 = 0 Step 3 : Solve y-2x2-3 = 0 Solving a Single Variable Equation : 3.1 Solve y-2x2-3 = 0

OpenStudy (anonymous):

thanks! but the only problem is my problem is a negative 2, and this one seems like it isn't?

OpenStudy (fibonaccichick666):

parentheses, they are important

OpenStudy (fibonaccichick666):

you don't need to solve anything, they just want you to plug and chug

OpenStudy (fibonaccichick666):

So what do you get for y if you let x=-3?

OpenStudy (anonymous):

I get 21, right? @FibonacciChick666